OpenStudy (anonymous):

find the product of (4x+3x)^2

OpenStudy (anonymous):

anyone plzz

OpenStudy (anonymous):

\[\left( 4x+3x \right)^{2}=\left( 4x \right)^{2}+\left( 3x \right)^{2}+2*4x*3x\] solve it. or \[\left( 4x+3x \right)^{2}=\left( 7x \right)^{2}=49x ^{2}\]

OpenStudy (anonymous):

okay thanks